### v2

# Underdetermined Fourier Extensions for Surface Partial Differential Equations

Abstract.
We analyze and test using Fourier extensions that minimize a Hilbert space norm for the purpose of solving partial differential equations (PDEs) on surfaces. In particular, we prove that the approach is arbitrarily high-order and also show a general result relating boundedness, solvability, and convergence that can be used to find eigenvalues. The method works by extending a solution to a surface PDE into a box-shaped domain so that the differential operators of the extended function agree with the surface differential operators, as in the Closest Point Method. This differs from approaches that require a basis for the surface of interest, which may not be available. Numerical experiments are also provided, demonstrating super-algebraic convergence. Current high-order methods for surface PDEs are often limited to a small class of surfaces or use radial basis functions (RBFs). Our approach offers certain advantages related to conditioning, generality, and ease of implementation. The method is meshfree and works on arbitrary surfaces (closed or non-closed) defined by point clouds with minimal conditions.
Key words and phrases:
numerical analysis; meshfree methods; Fourier extensions; spectral methods; partial differential equations; Hermite-Birkhoff interpolation, radial basis functions
1. Introduction
Partial differential equations (PDEs) on surfaces appear in a variety of contexts, often in medical imaging or computer graphics. Various methods for solving such problems have been developed over the past couple of decades; however, the number of methods with proofs of convergence is fairly limited.
Among existing approaches, surface finite element methods are likely the best understood, owing largely to the significant work of Dziuk and Elliott in the 2010s [dziuk13]. A number of implementations of radial basis function (RBF) methods, often using least squares collocation, have also been studied. RBFs are well understood for interpolation and can be high order, and while the popular Kansa’s [kansa90] is known to potentially fail, techniques based on oversampling have been successful on flat domains as well as surfaces when a point cloud regularity condition is satisfied [chen20]. The Closest Point Method (CPM) [ruuth08] has been extensively studied numerically, and consistency follows directly from consistency of the underlying interpolation and finite difference techniques, but stability is not yet fully understood, though the method works in practice in most cases. RBF methods using similar extension approaches as the CPM, either by performing a Closest Point extension directly [petra18], or by satisfying normal derivative conditions [piret12], have also been proposed.
We propose a new, flexible, and high-order class of meshfree methods for PDEs on surfaces using underdetermined Fourier extensions. These extensions are chosen to be norm-minimizing in a certain Hilbert space. We show that the norm-minimizing property is enough to prove statements regarding the methods’ convergence for a variety of problems. Finally, we connect our method to certain implementations of symmetric Hermite-Birkhoff interpolation with RBFs and demonstrate how it could be used to investigate the solvability of PDEs.
Norm-minimizing Hermite-Birkhoff interpolants have been previously studied in the RBF literature through Hermite RBFs [frank98, sun94]. More recently, Chandrasekaran, Gorman, and Mhaskar have studied solving the norm-minimization problem more directly [chand18], which has some numerical advantages. Namely, solving the optimization problem directly can improve conditioning and is often simpler to implement. In our work, we explore the direct solution of the norm-minimization problem for surface PDEs via function extension from surfaces to a box. Extension does not require a basis for functions on the surface; a standard Fourier basis can be used.
Along with numerical tests in Section 4, we present new analytical results in Propositions 9, 10, and 11 showing the high-order nature of our extension approach. We also prove a very general result regarding the relationship between boundedness, PDE solvability, and convergence in Proposition 12. In particular, we show that boundedness of numerical solutions to a PDE as point spacing goes to zero implies solvability of the non-discretized PDE and convergence of the numerical solution to a PDE solution, regardless of whether the problem has multiple solutions. This result is used to find eigenvalues in Subsection 4.2.
2. Preliminaries
2.1. Fourier Extensions
Spectral methods for PDEs typically work by searching for a PDE solution of the form:
$\tilde{u}=\sum_{n=1}^{N_{b}}a_{n}\phi_{n},$
where $\left\{a_{n}\right\}_{n=1}^{N_{b}}$ are coefficients, and $\left\{\phi_{n}\right\}_{n=1}^{N_{b}}$ are functions. $\left\{\phi_{n}\right\}_{n=1}^{N_{b}}$ are often chosen to be a subset of basis functions for some function space. Fourier basis functions are a common choice: $\phi_{n}\left(\boldsymbol{x}\right)=e^{i\boldsymbol{\omega}_{n}\cdot\boldsymbol{x}}$ for a set of frequencies $\left\{\boldsymbol{\omega}_{n}\right\}$. However, for all but the simplest domains, coming up with suitable functions $\left\{\phi_{n}\right\}$ can be challenging or impossible without an already existing, high-order numerical method.
A possible resolution to this problem is to avoid finding eigenfunctions on a complicated domain altogether. Instead, the possibly complicated domain $S$ is placed inside a box-shaped domain $\Omega\subseteq\mathbb{R}^{m}$. We can then use the box’s Fourier basis functions to expand functions on $S$, albeit not uniquely. In approximation theory terms, our functions $\left\{\phi_{n}\right\}_{n=1}^{N_{b}}$ will not form a basis for functions on $S$, but rather a frame [adcoc19].
Consider a reasonably nice function (where we will be more precise later) $f:S\to\mathbb{R}$. There exist many functions $\tilde{f}:\Omega\to\mathbb{R}$ such that
$f=\tilde{f}\biggr{|}_{S}.$
The idea of Fourier extension is to attempt to write a Fourier series for $\tilde{f}$:
$\tilde{f}=\sum_{n=1}^{\infty}a_{n}\phi_{n},$
where $\left\{a_{n}\right\}$ are coefficients and $\phi_{n}\left(\boldsymbol{x}\right)=e^{i\boldsymbol{\omega}_{n}\cdot\boldsymbol{x}}$ are Fourier basis functions for the box $\Omega$.
This sort of problem, where we extend $f$ from $S$ to a larger domain $\Omega\supset S$ using a Fourier series, is referred to by Boyd [boyd03] as Fourier extension of the third kind when $S$ and $\Omega$ have the same dimensionality. One motivation for computing such an extension is that we are able to compute derivatives of Fourier series on $\Omega$ with high accuracy.
Computing a Fourier extension without knowledge of the function’s exact Fourier coefficients is typically done by only approximately satisfying $f=\tilde{f}$ on $S$. A sample of points in $S$ is chosen, then the coefficients $a_{n}$ are computed via least squares. Certain implementations have been shown to converge super-algebraically in 1D for smooth functions when using Chebyshev nodes, or oversampled, uniformly spaced nodes [adcoc14].
Fourier extensions have been applied successfully in various contexts, ranging from straightforward 1D function approximation, to surface reconstruction [bruno07], to PDEs in flat domains [boyd05, stein17, bruno22, matth18]. Analytical results, however, remain fairly sparse in higher dimensions, particularly for PDEs.
2.2. Closest Point and Closest Point-like Extensions
Once we start working on manifolds embedded in a higher dimensional space, we have another motivation for performing extensions from a manifold $S$ to a larger domain $\Omega$ of co-dimension zero. There are many available methods for computing derivatives and solving PDEs in flat, Euclidean spaces; conversely, solving PDEs on manifolds directly can be quite challenging, especially if a connected mesh is not available.
A straightforward method for extending functions off of manifolds is a Closest Point extension, which is the idea underlying the Closest Point Method (CPM) for PDEs on surfaces [ruuth08].
Definition 1.
The (Euclidean) Closest Point extension of a function $f:S\to\mathbb{R}$, is the function $f\circ\text{cp}$, where
$\text{cp}\left(\boldsymbol{x}\right)=\underset{\boldsymbol{y}\in S}{\text{argmin}}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|_{2}.$
If $S$ is a twice differentiable manifold, then cp is well-defined and differentiable in a neighbourhood of an open subset of $S$. The utility of such an extension is clear from the following facts (see [ruuth08] for the original statements and [marz12] for detailed proofs).
Fact 1.
If $f\in C^{1}\left(S\right)$ where $S$ is a twice continuously differentiable manifold, then
$\nabla\left(f\circ\text{cp}\right)\biggr{|}_{S}=\nabla_{S}f,$
where $\nabla_{S}$ is the gradient intrinsic to $S$.
Fact 2.
If $f\in C^{2}\left(S\right)$ where $S$ is a thrice differentiable manifold, then
$\Delta\left(f\circ\text{cp}\right)\biggr{|}_{S}=\Delta_{S}f,$
where $\Delta_{S}$ is the Laplace-Beltrami operator on $S$.
Combining these facts allows for various PDEs to be solved on surfaces without a parametrization or a mesh.
As it turns out, Closest Point extensions are not the only extensions where Propositions 1 and 2 hold. März and Macdonald studied extensions of the form $f\circ P$, where $P:\Omega\to S$ is idempotent, and found general conditions needed for $P$ so that Propositions 1 and 2 still hold with $P$ instead of cp [marz12].
We state a couple of facts needed for functions from $\Omega\to S$ in general to be used for computing surface differential operators. First, we need a definition.
Definition 2.
Let $f\in C^{1}\left(S\right)$ where $S$ is a twice differentiable manifold. Then $\tilde{f}\in C^{1}\left(\Omega\right)$ is a first-order Closest Point-like extension of $f$ when
$\hat{\boldsymbol{n}}_{S}\cdot\nabla\tilde{f}\biggr{|}_{S}=0\quad\text{and}\quad\tilde{f}\biggr{|}_{S}=f,$
where $\hat{\boldsymbol{n}}_{S}$ is the normal vector to $S$.
This definition is motivated by the next fact.
Proposition 1.
Let $f\in C^{1}\left(S\right)$ and $\tilde{f}\in C^{1}\left(\Omega\right)$ be a first-order Closest Point-like extension of $f$. Then
$\nabla\tilde{f}\biggr{|}_{S}=\nabla_{S}f.$
To work with the Laplace-Beltrami operator, we need an important result adapted from Xu and Zhao [xu03].
Lemma 2.
[xu03, Lemma 1] Let $f\in C^{2}\left(S\right),\tilde{f}\in C^{2}\left(\Omega\right)$ where $S$ is a thrice differentiable manifold. If
$\tilde{f}\biggr{|}_{S}=f,$
then
$\Delta_{S}f=\left(\Delta\tilde{f}-\kappa\hat{\boldsymbol{n}}_{S}\cdot\nabla\tilde{f}-\hat{\boldsymbol{n}}_{S}\cdot\left(D^{2}\tilde{f}\right)\hat{\boldsymbol{n}}_{S}\right)\biggr{|}_{S},$
where $\kappa=\nabla_{S}\cdot\hat{\boldsymbol{n}}_{S}$ is the mean curvature (sum of principal curvatures in our convention) of $S$.
Using Lemma 2, we can then compute $\Delta_{S}$ using a first-order Closest Point-like extension.
Corollary 3.
Let $f\in C^{2}\left(S\right)$ and let $\tilde{f}\in C^{2}\left(\Omega\right)$ be a first-order Closest Point-like extension of $f$. Then
$\nabla\tilde{f}\biggr{|}_{S}=\nabla_{S}f,\quad\left(\Delta\tilde{f}-\hat{\boldsymbol{n}}_{S}\cdot\left(D^{2}\tilde{f}\right)\hat{\boldsymbol{n}}_{S}\right)\biggr{|}_{S}=\Delta_{S}f.$
Crucially, the only surface information we need is $\hat{\boldsymbol{n}}_{S}$, not any of its derivatives or curvature information. Also note that Corollary 3 only imposes one additional interpolation condition. Some previous methods, such as the higher-order version of Piret’s orthogonal gradients method with RBFs [piret12], impose two additional conditions; $\hat{\boldsymbol{n}}_{s}\cdot\nabla_{S}f$ and $\hat{\boldsymbol{n}}_{S}\cdot\left(D^{2}\tilde{f}\right)\hat{\boldsymbol{n}}_{S}$ are both set to zero. Using only one condition is more computationally efficient. We use this corollary to solve surface PDEs in Section 4.
We now introduce one more fact and a definition before another preliminary result.
Proposition 4.
For every $C^{p}$ manifold $S\subset\mathbb{R}^{n}$, for small enough $\beta>0$, where $S\subset U_{\beta}$ and $U_{\beta}$ contains all points with distance less than $\beta$ to $S$, there exists a signed distance function $\varphi_{\beta}\in C^{p}\left(U_{\beta}\right)$. That is, $\varphi_{\beta}$ is $C^{p}$ and $\left|\varphi_{\beta}\left(\boldsymbol{x}\right)\right|=\inf_{\boldsymbol{y}\in S}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|_{2}<\beta$.
Proposition 4 is a simple consequence of Lemma 14.16 from [gilba01].
Definition 3.
A $C^{\infty}$ top hat function is a function $\Phi\in C^{\infty}\left(\mathbb{R}\right)$ such that for some $\alpha,\beta>0$, where $\alpha<\beta$,
$\Phi\left(x\right)$ $=1\text{, for $x\in\left[-\alpha,\alpha\right]$}$
$\Phi\left(x\right)$ $=0\text{, for $x\notin\left[-\beta,\beta\right]$}.$
There are many examples of smooth top hat functions [boyd06]. Using the Euclidean Closest Point function and top hat functions, we can now show that smooth, periodic, Closest Point-like extensions exist.
Proposition 5.
For $p\geq 1$, $f\in C^{p}\left(S\right)$, and a $C^{p+1}$ manifold $S$, there exists a periodic function $\tilde{f}\in C^{p}\left(\Omega\right)$ such that $\tilde{f}$ is a first-order Closest Point-like extension of $f$.
This is constructed as $\left(f\circ\text{cp}\right)\cdot\left(\Phi\circ\varphi_{\beta}\right)$ within a distance $\beta$ of $S$ and zero elsewhere, where $\varphi_{\beta}$ is as in Proposition 4 and $\Phi$ is a smooth top hat function with compact support of width $2\beta$ as in Definition 3. Note that $\text{cp}\left(\boldsymbol{x}\right)=\boldsymbol{x}-\varphi_{\beta}\left(\boldsymbol{x}\right)\nabla\varphi_{\beta}\left(\boldsymbol{x}\right)$, which is why $S$ must be $C^{p+1}$. This result tells us that for the PDEs on surfaces we consider, there are periodic extended solutions with the same degree of smoothness as the solution on the surface. In particular, the constrained norm-minimization problems that we set up will have feasible solutions.
2.3. Radial Basis Function Interpolation
We will make use of existing interpolation results to prove the convergence of our own methods; we therefore review some basic results regarding RBF interpolation.
Radial basis functions (RBFs) are a class of functions used for interpolation and numerical PDEs that depend on the location of the interpolation points $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}\subset S\subset\mathbb{R}^{n}$. This is vital since, due to the Mairhuber-Curtis Theorem [curti59, mairh56], any set of $\tilde{N}$ functions that does not depend on the location of interpolation points cannot uniquely interpolate every set of $\tilde{N}$ points. Specifically, polynomial interpolation when the number of polynomials matches the number of points, can fail to produce a solution.
Before stating some results, we first define the fill distance.
Definition 4.
Let $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}\subset S\subset\mathbb{R}^{m}$ be distinct points. Define the fill distance of $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}$ in $S$:
$h_{\text{max}}:=\sup_{\boldsymbol{x}\in S}\min_{k\in\left\{1,2,\ldots,\tilde{N}\right\}}\left\|\boldsymbol{x}-\boldsymbol{x}_{k}\right\|_{2}.$
In 1D, with $S=\left[a,b\right]$ and $\left\{x_{k}\right\}_{1}^{\tilde{N}}$ increasing, this is equivalent to
$h_{\text{max}}=\max\left(\left\{\frac{1}{2}\left|x_{k}-x_{k-1}\right|\right\}_{k=2}^{\tilde{N}}\cup\left\{x_{1}-a,b-x_{\tilde{N}}\right\}\right).$
That is, the fill distance is the largest distance between a point $\boldsymbol{x}\in S$ and its closest point in $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}$.
As suggested by the name, RBFs are radially symmetric. We also restrict ourselves to the discussion of positive definite RBFs, which we now define.
Definition 5.
A positive definite radial basis function is a function $\phi:\mathbb{R}^{m}\to\mathbb{R}$ that posesses two properties:
  1. (1)
    $\phi$ is radial;
    $\left\|\boldsymbol{x}\right\|_{2}=\left\|\boldsymbol{y}\right\|_{2}\implies\phi\left(\boldsymbol{x}\right)=\phi\left(\boldsymbol{y}\right).$
  2. (2)
    $\phi$ is (strictly) positive definite. That is, if we define the interpolation matrix $\boldsymbol{\Phi}$ on any set of unique points $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}\subset\mathbb{R}^{m}$:
    $\boldsymbol{\Phi}_{jk}=\phi\left(\boldsymbol{x}_{j}-\boldsymbol{x}_{k}\right),$
    then $\boldsymbol{\Phi}$ is a symmetric positive definite matrix.
Note that this means linear combinations of the functions $\left\{\psi_{k}\right\}_{k=1}^{\tilde{N}}$ defined by $\psi_{k}\left(\boldsymbol{x}\right)=\phi\left(\boldsymbol{x}-\boldsymbol{x}_{k}\right)$ can interpolate any function on $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}$ (since $\boldsymbol{\Phi}$ must be positive definite). Given a function $f$ and RBF $\phi$, we refer to the unique function $\tilde{f}\in\text{span}\left\{\psi_{k}\right\}$ such that $\tilde{f}\left(\boldsymbol{x}_{k}\right)=f\left(\boldsymbol{x}_{k}\right)$ for each $k\in\left\{1,2,\ldots,\tilde{N}\right\}$ as the RBF interpolant of $f$ on $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}$ using $\phi$. Our goal now is to quickly prove a general statement regarding functions with scattered zeros.
First, we need a definition (see Definition 3.6 of [wendl04]) to restrict the types of domains we can consider.
Definition 6.
A set $\Omega\subset\mathbb{R}^{m}$ satisfies an interior cone condition if there exists an angle $\theta\in\left(0,\frac{\pi}{2}\right)$ and a radius $r>0$ such that for all $\boldsymbol{x}\in\Omega$, there exists some unit vector $\boldsymbol{\xi}\left(\boldsymbol{x}\right)$ so that
$\left\{\boldsymbol{x}+\lambda\boldsymbol{y}:\boldsymbol{y}\in\mathbb{R}^{d},\left\|\boldsymbol{y}\right\|_{2}=1,\boldsymbol{y}\cdot\boldsymbol{\xi}\left(\boldsymbol{x}\right)\geq\cos\theta,\lambda\in\left[0,r\right]\right\}\subseteq\Omega.$
In other words, $\Omega$ satisfies an interior cone condition if there is some cone of a fixed size and angle that can be placed with its vertex at each $\boldsymbol{x}\in\Omega$ and remain entirely contained in $\Omega$. Informally, $\Omega$ is only “finitely pointy”.
The next proposition is adapted from Wendland [wendl04].
Proposition 6.
[wendl04, Thms. 10.35 and 11.17] Let $f\in H^{\frac{m}{2}+q+\frac{1}{2}}\left(\tilde{\Omega}\right)$ with $m\geq 3$ if $q=0$ and $m\in\mathbb{N}$ for $q>0$ where $\tilde{\Omega}\supset\Omega$ is bounded and $\Omega\subset\mathbb{R}^{m}$ satisfies an interior cone condition. Then there exists a radial basis function $\phi_{m,q}$ such that if $\tilde{f}$ is the RBF interpolant of $f$ on $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}$ using $\phi_{m,q}$ where $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}$ has fill distance $h_{\text{max}}$ on $\Omega$, then there exist constants $C_{m,q},h_{0,m,q}>0$ such that as long as $h_{\text{max}}\leq h_{0,m,q}$,
$\left\|f-\tilde{f}\right\|_{L^{\infty}\left(\Omega\right)}\leq C_{m,q}h_{\text{max}}^{q+\frac{1}{2}}\left\|f\right\|_{H^{\frac{m}{2}+q+\frac{1}{2}}\left(\Omega\right)}.$
Noting that the RBF interpolant of a function that vanishes on $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}$ is simply the zero function, we then have a general result for functions vanishing on scattered points.
Corollary 7.
Let $f\in H^{\frac{m}{2}+q+\frac{1}{2}}\left(\tilde{\Omega}\right)$ with $m\geq 3$ if $q=0$ and $m\in\mathbb{N}$ for $q>0$ where $\tilde{\Omega}\supset\Omega$ is bounded and $\Omega$ satisfies an interior cone condition. Assume $f$ vanishes on $\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}$ with fill distance $h_{\text{max}}$ on $\Omega$, then for all $p\in\left\{0,1,\ldots q\right\}$ for $m\geq 3$ and $p\in\left\{1,\ldots q\right\}$ for $m<3$ there exist constants $C_{m,q},h_{0,m,q}>0$ such that as long as $h_{\text{max}}\leq h_{0,m,q}$,
$\left\|f\right\|_{L^{\infty}\left(\Omega\right)}\leq C_{m,q}h_{\text{max}}^{p+\frac{1}{2}}\left\|f\right\|_{H^{\frac{m}{2}+p+\frac{1}{2}}\left(\Omega\right)}.$
Similar results regarding error bounds for functions with scattered zeros exist in the literature, and could also be used for efficient convergence proofs for a wide variety of methods, including our own. In particular, more general statements for convergence in Sobolev norms have been shown by Narcowich, Ward, and Wendland [narco05].
2.3.1. Application to Other Interpolation Methods
Corollary 7 has implications for other interpolation methods, noting that it is a general result; it does not have any specific relation to RBF interpolation other than the method of proof. This gives an important insight into constructing effective interpolation methods. If we are able to bound various $H^{k}$ norms of our error $u-\tilde{u}$, where $\tilde{u}$ is an interpolant and $u$ is the original function, then Corollary 7 guarantees uniform, high-order convergence of $\tilde{u}$ to $u$ as the fill distance goes to zero.
We also note that a version of Corollary 7 holds on manifolds $S$ as well, if the manifold is sufficiently smooth. While the manifold itself may not satisfy an interior cone condition, the domain $U$ of a local parametrization $\sigma:U\to S$ typically will. The corollary can then be applied to $f\circ\sigma$.
3. Proposed Methods & Analysis
In order to solve PDEs on scattered data (such as a point cloud on a surface), we need to be able to interpolate functions over scattered data on arbitrary domains. RBFs, as discussed in Subsection 2.3, provide one way of doing this.
Among existing approaches are interpolation method based on using a standard RBF basis, such as the Piret’s method [piret12]. Like their flat domain counterparts such as Kansa’s original method [kansa90], these methods are often quite successful in practice, but face some theoretical issues due to the possible singularity of the interpolation matrix once derivatives are involved. As with most RBF methods, the accuracy of these methods is often severely limited by conditioning, leading to errors much higher than machine error.
Other existing approaches are RBF-FD methods, which are particularly useful for large, time-dependent problems due to the sparsity of the differentiation matrices produced. These methods can either work combined with the Closest Point Method [petra18] or by more directly approximating the surface operators using estimates of a local level set [alvar21]. Being local methods, such methods typically have limited orders of convergence. Furthermore, stability (for time-stepping) and non-singularity (for elliptic problems) of such methods is currently unknown.
Lastly, various implementations using standard RBF basis functions and least-squares collocation have also been used. Significant analysis of these sorts of approaches, along with numerical testing, has been done by Chen and Ling [chen20]. A quasi-uniform point cloud is required, and a possibly quite dense point cloud of collocation points is required relative to density of the point cloud of RBF centres for guaranteed convergence. Alternatively, one may approximate a weak form of the PDE to achieve symmetry of the discretized Laplace-Beltrami operator, as in the work by Yan, Jiang, and Harlim [yan23], which overcomes a weakness of standard interpolation approaches, but requires knowledge of the distribution from which collocation points are sampled, which may not be available.
We seek to explore other interpolation methods based on underdetermined Fourier extensions, which will allow for super-algebraic convergence with minimal conditions on the point clouds that can be used. However, as we will see in Subsection 3.3, our methods can also be viewed as a sort of periodic Hermite-Birkhoff RBF interpolation with some numerical advantages.
Existing Fourier extension methods typically use an overdetermined system. Such methods for function approximation and bulk PDE problems in 2D have been recently proposed [stein17, matth18, bruno22], though convergence analysis is typically limited to 1D. Analysis of overdetermined methods also relies on relating the discrete least squares problem to an $L^{2}$-norm minimization problem through quadrature. Such an approach would not be feasible on all but a few surfaces where high-order quadrature schemes can be developed.
3.1. Hermite-Birkhoff Fourier Extension Problem
We first set up a constrained optimization problem. Let $S\subset\mathbb{R}^{m}$ be our domain of interest and let $\Omega\supset S$ be a box. Select a point cloud $S_{\tilde{N}}:=\left\{\boldsymbol{x}_{k}\right\}_{k=1}^{\tilde{N}}\subset S$ and linear differential operators $\left\{\mathcal{F}_{k}\right\}_{k=1}^{\tilde{N}}$ of maximum order $p\in\mathbb{N}$, where $\mathcal{F}_{k}=\sum_{\left|\alpha\right|\leq p}c_{k,\alpha}\partial^{\alpha}$ and the coefficient functions $c_{k,\alpha}$ are bounded and defined in a neighbourhood of $\boldsymbol{x}_{k}$. Let $\left\{\phi_{n}\right\}_{n=1}^{\infty}$ be the Fourier basis for the box consisting of complex exponential eigenfunctions of the Lapacian on $\Omega$ so that $\phi_{n}\left(\boldsymbol{x}\right)=e^{i\boldsymbol{\omega}_{n}\cdot\boldsymbol{x}}$ for some frequency $\boldsymbol{\omega}_{n}$. Let $d>0$ be a sequence and let $\left\{f_{k}\right\}_{k=1}^{\tilde{N}}\subset\mathbb{C}$ be interpolation values. We then consider the problem:
(3.1) minimize, over $b\in\ell^{2}$: $\left\|b\right\|_{\ell^{2}}$
subject to: $\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}b_{n}\left(\mathcal{F}_{k}\phi_{n}\right)\left(\boldsymbol{x}_{k}\right)=f_{k}\text{, for $k\in\left\{1,2,\ldots,\tilde{N}\right\}$}.$
As long as the Hermite-Birkhoff interpolation conditions at each distinct point are consistent, there will be at least one feasible solution to this problem. In fact, there is a feasible sequence $b$ with a finite number of non-zero terms (see Remark 2 in Subsection 3.6 for a simple upper bound on the minimum number of terms needed). We now show uniqueness.
Proposition 8.
If the feasible set for (3.1) is non-empty and $d^{-\frac{1}{2}}\left\|\boldsymbol{\omega}\right\|_{2}^{p}\in\ell^{2}$, where we use the notation $\left\|\boldsymbol{\omega}\right\|_{2}^{p}$ to mean the sequence $\left\{\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{p}\right\}_{n=1}^{\infty}$, then the solution to (3.1) is unique.
Proof.
Define the linear operator $\mathcal{L}_{\text{HB}}:\ell^{2}\to\mathbb{C}^{\tilde{N}}$ by:
$\left(\mathcal{L}_{\text{HB}}b\right)_{k}:=\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}b_{n}\left(\mathcal{F}_{k}\phi_{n}\right)\left(\boldsymbol{x}_{k}\right).$
If $d^{-\frac{1}{2}}\left\|\boldsymbol{\omega}\right\|_{2}^{p}\in\ell^{2}$, then noting that each $\mathcal{F}_{k}$ is of order at most $p$ and has bounded coefficient functions, it is clear that $\mathcal{L}_{\text{HB}}$ is bounded, since for $q\in\left\{0,1,\ldots,p\right\}$,
[…]
Since $\mathcal{L}_{\text{HB}}$ is bounded, $\mathcal{N}\left(\mathcal{L}_{\text{HB}}\right)$ is closed and the constraint set is closed and convex. By a standard result in functional analysis (see, for example, 3.3-1 of Kreyszig [kreys91]), if the constraint set is non-empty, there is a unique element with a minimum norm; this is the unique solution. ∎
3.2. Native Hilbert Spaces
Let $\tilde{b}$ be the solution to (3.1), then our chosen Hermite-Birkhoff interpolant is
(3.2) $\tilde{u}=\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}\tilde{b}_{n}\phi_{n}.$
Our chosen solution $\tilde{u}$ turns out to be an element of a particular Hilbert space of functions constructed from an isometry to $\ell^{2}$.
Definition 7.
Let $\left\{\phi_{n}\right\}_{n=1}^{\infty}$ be the Fourier basis for the box $\Omega$ with periodic boundary conditions ($\phi_{n}\left(\boldsymbol{x}\right)=e^{i\boldsymbol{\omega}_{n}\cdot\boldsymbol{x}}$), then we define the Hilbert space $\mathcal{H}\left(d\right)$ by
$\mathcal{H}\left(d\right):=\left\{\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}b_{n}\phi_{n}:b\in\ell^{2}\right\}.$
The associated inner product for $\mathcal{H}\left(d\right)$ is
$\left(\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}a_{n}\phi_{n},\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}b_{n}\phi_{n}\right)_{d}=\left(a,b\right)_{\ell_{2}}=\sum_{n=1}^{\infty}d_{n}\left(d_{n}^{-\frac{1}{2}}a_{n}\right)\left(d_{n}^{-\frac{1}{2}}b_{n}\right)^{*}.$
Alternatively, if $u,v\in\mathcal{H}\left(d\right)$ have sequences of Fourier coefficients $\hat{u},\hat{v}$, respectively, then
$\left(u,v\right)_{d}$ $=\sum_{n=1}^{\infty}d_{n}\hat{u}_{n}\hat{v}_{n}^{*},\quad\left\|u\right\|_{d}=\sqrt{\sum_{n=1}^{\infty}d_{n}\left|\hat{u}_{n}\right|^{2}}=\left\|d^{\frac{1}{2}}\hat{u}\right\|_{\ell^{2}}.$
For example, if $d=\left(1+\left\|\boldsymbol{\omega}\right\|_{2}^{2}\right)^{q}$, then $\mathcal{H}\left(d\right)=H^{q}\left(\Omega\right)$; this would be a form of minimum Sobolev norm interpolation [chand13, chand18], which has been used for PDEs in flat domains previously in a different setup [chand15]. Note that we often refer to $\left\|\cdot\right\|_{d}$ as the $d$-norm.
Remark 1.
$\tilde{u}$ is the solution to the constrained optimization problem:
minimize, over $u\in\mathcal{H}\left(d\right)$: $\left\|u\right\|_{d}$
subject to: $\left(\mathcal{F}_{k}u\right)\left(\boldsymbol{x}_{k}\right)=f_{k}\text{, for each }k\in\left\{1,2,\ldots,\tilde{N}\right\}.$
3.3. The Adjoint Range and Connection to Radial Basis Functions
Recall that we want to match certain derivative conditions at each point $\boldsymbol{x}_{k}\in S$; there is some collection of linear differential operators $\left\{\mathcal{F}_{k}\right\}_{k=1}^{\tilde{N}}$ so that
$\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}\tilde{b}_{n}\left(\mathcal{F}_{k}\phi_{n}\right)\left(\boldsymbol{x}_{k}\right)=f_{k}\text{, for each $k\in\left\{1,2,\ldots,\tilde{N}\right\}$},$
where $\tilde{b}$ is again the solution to (3.1). Then, where $p$ is the maximum order of $\mathcal{F}_{k}$ over all $k\in\left\{1,2,\ldots,\tilde{N}\right\}$, $\mathcal{L}_{\text{HB}}:\ell^{2}\to\mathbb{C}^{\tilde{N}}$ is bounded as long as $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$ (see Proposition 8), and elementary functional analysis gives us:
$\mathcal{L}_{\text{HB}}$ $:b\mapsto\left(\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}\left(\mathcal{F}_{k}\phi_{n}\right)\left(\boldsymbol{x}_{k}\right)b_{n}\right)_{k=1}^{\tilde{N}}$
$\mathcal{L}_{\text{HB}}^{*}$ $:\boldsymbol{\beta}\mapsto\left(\sum_{k=1}^{\tilde{N}}d_{n}^{-\frac{1}{2}}\left(\mathcal{F}_{k}\phi_{n}\right)^{*}\left(\boldsymbol{x}_{k}\right)\beta_{k}\right)_{n=1}^{\infty}$
(3.3) $\tilde{b}$ $\in\mathcal{N}\left(\mathcal{L}_{\text{HB}}\right)^{\perp}=\mathcal{R}\left(\mathcal{L}_{\text{HB}}^{*}\right)=\text{span}\left\{\left(d_{n}^{-\frac{1}{2}}\left(\mathcal{F}_{k}\phi_{n}\right)^{*}\left(\boldsymbol{x}_{k}\right)\right)_{n=1}^{\infty}\right\}_{k=1}^{\tilde{N}}.$
We then define a set of functions $\left\{\psi_{k}\right\}_{1}^{\tilde{N}}$ by:
$\psi_{k}$ $:=\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}\left(d_{n}^{-\frac{1}{2}}\left(\mathcal{F}_{k}\phi_{n}\right)^{*}\left(\boldsymbol{x}_{k}\right)\right)\phi_{n}.$
Then, there exists some function $\tilde{u}\in\text{span}\left\{\psi_{k}\right\}_{k=1}^{\tilde{N}}$ so that for each $i$,
$\left(\mathcal{F}_{i}\tilde{u}\right)\left(\boldsymbol{x}_{i}\right)=f_{i}.$
Furthermore, the Fourier coefficients of $\tilde{u}$ minimize $\sum_{n=1}^{\infty}d_{n}\left|a_{n}\right|^{2}$ over all possible choices of Fourier coefficients $a_{n}$ that match the constraint. Finally, let: $\tilde{u}=\sum_{j=1}^{\tilde{N}}\beta_{j}\psi_{j}$, then we want $\boldsymbol{\beta}=\left(\beta_{j}\right)_{j=1}^{\tilde{N}}\in\mathbb{C}^{\tilde{N}}$ to solve
$\sum_{j=1}^{\tilde{N}}\beta_{j}\left(\sum_{n=1}^{\infty}d_{n}^{-1}\left(\mathcal{F}_{j}\phi_{n}\right)^{*}\left(\boldsymbol{x}_{j}\right)\left(\mathcal{F}_{i}\phi_{n}\right)\left(\boldsymbol{x}_{i}\right)\right)=\sum_{j=1}^{\tilde{N}}\boldsymbol{\Phi}_{ij}\beta_{j}=f_{i},$
where
$\boldsymbol{\Phi}_{ij}=\sum_{n=1}^{\infty}d_{n}^{-1}\left(\mathcal{F}_{j}\phi_{n}\right)^{*}\left(\boldsymbol{x}_{j}\right)\left(\mathcal{F}_{i}\phi_{n}\right)\left(\boldsymbol{x}_{i}\right)=\boldsymbol{\Phi}_{ji}^{*}.$
Therefore, the system that must be solved for $\boldsymbol{\beta}$ is self-adjoint. Also, for any $\boldsymbol{\alpha}\in\mathbb{C}^{\tilde{N}}$,
$\boldsymbol{\alpha}^{*}\boldsymbol{\Phi}\boldsymbol{\alpha}$ $=\sum_{i=1}^{\tilde{N}}\sum_{j=1}^{\tilde{N}}\left(\sum_{n=1}^{\infty}d_{n}^{-1}\left(\mathcal{F}_{j}\phi_{n}\right)^{*}\left(\boldsymbol{x}_{j}\right)\left(\mathcal{F}_{i}\phi_{n}\right)\left(\boldsymbol{x}_{i}\right)\right)\alpha_{i}^{*}\alpha_{j}$
$=\sum_{n=1}^{\infty}\left(\sum_{i=1}^{\tilde{N}}d_{n}^{-\frac{1}{2}}\left(\mathcal{F}_{i}\phi_{n}\right)\left(\boldsymbol{x}_{i}\right)\alpha_{i}^{*}\right)\left(\sum_{j=1}^{\tilde{N}}d_{n}^{-\frac{1}{2}}\left(\mathcal{F}_{j}\phi_{n}\right)^{*}\left(\boldsymbol{x}_{j}\right)\alpha_{j}\right)$
$=\sum_{n=1}^{\infty}\left(\mathcal{L}_{\text{HB}}^{*}\boldsymbol{\alpha}\right)_{n}^{*}\left(\mathcal{L}_{\text{HB}}^{*}\boldsymbol{\alpha}\right)_{n}=\left\|\mathcal{L}_{\text{HB}}^{*}\boldsymbol{\alpha}\right\|_{\ell^{2}}^{2}\geq 0.$
Thus, $\boldsymbol{\Phi}$ is positive semi-definite. Moreover, $\boldsymbol{\alpha}^{*}\boldsymbol{\Phi}\boldsymbol{\alpha}\neq 0$ due to the existence of an interpolant (namely, the norm-minimizing interpolant in $\mathcal{H}\left(d\right)$) in $\text{span}\left\{\psi_{k}\right\}_{k=1}^{\tilde{N}}$ for any set of interpolation values $\left\{f_{i}\right\}_{i=1}^{\tilde{N}}$. This implies uniqueness of $\boldsymbol{\beta}$ since the system is square. That is, for any set of interpolation conditions $\left(\mathcal{F}_{i}\tilde{u}\right)\left(\boldsymbol{x}_{i}\right)=f_{i}$ for $i\in\left\{1,2,\ldots,\tilde{N}\right\}$, there is a unique, $d$-norm minimizing interpolant in $\text{span}\left\{\psi_{k}\right\}_{k=1}^{\tilde{N}}$ such that the system to find the coefficients of the interpolant in the $\left\{\psi_{k}\right\}_{k=1}^{\tilde{N}}$ basis is self-adjoint and positive definite.
Also note that $\left(\psi_{k},\psi_{j}\right)_{d}=\left(\mathcal{L}_{\text{HB}}^{*}\boldsymbol{e}_{k},\mathcal{L}_{\text{HB}}^{*}\boldsymbol{e}_{j}\right)_{\ell^{2}}=\boldsymbol{\Phi}_{kj}$. This is closely related to symmetric Hermite RBF methods for PDEs [sun94, frank98], which have been applied to problems in flat domains and rely on a similar norm-minimizing property.
3.4. Setting up a PDE
We now set up a PDE problem. Let $\Omega^{\left(j\right)}\subseteq\Omega$ for $j\in\left\{1,2,\ldots,N_{\Omega}\right\}$ all have co-dimension zero and let $S^{\left(j\right)}\subset\Omega$ for $j\in\left\{1,2,\ldots,N_{S}\right\}$ all be manifolds of co-dimension greater than zero or single points, where $\Omega$ is a box-shaped domain in $\mathbb{R}^{m}$. Define point clouds $\Omega_{\tilde{N}_{j}}^{\left(j\right)}:=\left\{\boldsymbol{x}_{k}^{\left(j\right)}\right\}_{k=1}^{\tilde{N}_{j}}\subset\Omega^{\left(j\right)}$ and $S_{\tilde{M}_{j}}^{\left(j\right)}:=\left\{\boldsymbol{y}_{k}^{\left(j\right)}\right\}_{k=1}^{\tilde{M}_{j}}\subseteq S^{\left(j\right)}$. We consider the problem:
(3.4) $\mathcal{F}^{\left(j\right)}u\biggr{|}_{\Omega^{\left(j\right)}}$ $=f^{\left(j\right)}\text{, for each $j\in\left\{1,2,\ldots,N_{\Omega}\right\}$}$
$\mathcal{G}^{\left(j\right)}u\biggr{|}_{S^{\left(j\right)}}$ $=g^{\left(j\right)}\biggr{|}_{S^{\left(j\right)}}\text{, for each$j\in\left\{1,2,\ldots,N_{S}\right\}$},$
where $\mathcal{F}^{\left(j\right)}$ and $\mathcal{G}^{\left(j\right)}$ are linear differential operators of order at most $p$, $f^{\left(j\right)}$ is a function defined on $\Omega^{\left(j\right)}$, and $g^{\left(j\right)}$ is defined in a neighbourhood of $S^{\left(j\right)}$ for each $j$. Note that we can extend functions on $S^{\left(j\right)}$ to an open neighbourhood of $S^{\left(j\right)}$ while preserving their differentiability using Proposition 5, as long as $S^{\left(j\right)}$ is $C^{2}$; we note this here since the degree of differentiability of $g^{\left(j\right)}$ shows up in convergence rates in Subsection 3.5. This form covers all strong-form linear PDEs on piecewise smooth manifolds we may consider. We then have the discretization:
(3.5) minimize, over $b\in\ell^{2}$: $\left\|b\right\|_{\ell^{2}}$
subject to: $\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}b_{n}\left(\mathcal{F}^{\left(j\right)}\phi_{n}\right)\left(\boldsymbol{x}_{k}^{\left(j\right)}\right)=f^{\left(j\right)}\left(\boldsymbol{x}_{k}^{\left(j\right)}\right)\text{, foreach $j,k$}$
$\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}b_{n}\left(\mathcal{G}^{\left(j\right)}\phi_{n}\right)\left(\boldsymbol{y}_{k}^{\left(j\right)}\right)=g^{\left(j\right)}\left(\boldsymbol{y}_{k}^{\left(j\right)}\right)\text{, foreach $j,k$}.$
Assuming $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$, let $\tilde{u}=\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}\tilde{b}_{n}\phi_{n}$ as before, where $\tilde{b}$ is the unique solution to (3.5), and the functions $\left\{\phi_{n}\right\}$ are again Fourier modes on the box $\Omega$ so that $\phi_{n}\left(\boldsymbol{x}\right)=e^{i\boldsymbol{\omega}_{n}\cdot\boldsymbol{x}}$. If $u$ is a solution to the original PDE (3.4), extended to $\Omega$, and $u\in\mathcal{H}\left(d\right)$, then $u$ is a feasible Hermite-Birkhoff interpolant and $\left\|\tilde{u}\right\|_{d}\leq\left\|u\right\|_{d}$. Note that we can always choose $d$ to satisfy $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$ to be guaranteed a unique solution to (3.5) and have $\left\|\tilde{u}\right\|_{d}\leq\left\|u\right\|_{d}<\infty$ as long as an exact solution $u$ is sufficiently smooth. For certain surface PDEs involving the Laplace-Beltrami operator with sufficiently smooth solutions on the surface, an extended, periodic solution $u$ on $\Omega$ with the same number of continuous derivatives will exist due to Proposition 5.
3.5. Analysis for Hermite-Birkhoff Problem
Let $h_{\text{max}}$ be the largest fill distance of each point cloud $\Omega_{\tilde{N}_{j}}^{\left(j\right)}$ and $S_{\tilde{M}_{j}}^{\left(j\right)}$ in their respective domains $\Omega^{\left(j\right)}$ and $S^{\left(j\right)}$. We can then prove uniform convergence of $\mathcal{F}^{\left(j\right)}\tilde{u}\to f^{\left(j\right)}$ and $\mathcal{G}^{\left(j\right)}\tilde{u}\to g^{\left(j\right)}$; convergence is high-order or super-algebraic with respect to $h_{\text{max}}$ as long as certain smoothness requirements are met. This, or similar arguments with different norms, is enough to prove high-order convergence of $\tilde{u}$ to a solution to the PDE if the PDE is well-posed with respect to the norms in consideration. Proposition 12 will later show convergence to a solution of the PDE directly under weaker conditions, albeit without rate estimates.
Note for the following results that we typically omit the notation for the restriction of a function when the restricted domain is otherwise clear, such as in a norm. For example, we may write $\left\|\tilde{u}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}$ to mean $\left\|\tilde{u}|_{S^{\left(j\right)}}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}$. We also note again here that we use $\left\|\boldsymbol{\omega}\right\|_{2}$ to mean the sequence $\left\{\left\|\boldsymbol{\omega}_{n}\right\|_{2}\right\}_{n=1}^{\infty}$ (it is a sequence of finite dimensional 2-norms, not an $\ell^{2}$ norm).
Proposition 9.
Let $\tilde{b}$ be the solution to (3.5) and let $\tilde{u}=\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}\tilde{b}_{n}\phi_{n}$. Assume $u\in\mathcal{H}\left(d\right)$ is any solution to (3.4). Let $j\in\left\{1,2,\ldots,N_{\Omega}\right\}$ be fixed. Assume $\Omega^{\left(j\right)}$ satisfies an interior cone condition, and assume $q\in\mathbb{N}$ is such that $f^{\left(j\right)}\in H^{q}\left(S\right)$ and $F^{\left(j,q\right)}d^{-\frac{1}{2}}\in\ell^{2}$, where $F^{\left(j,q\right)}:=\left\{\left\|\mathcal{F}^{\left(j\right)}\phi_{n}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}\right\}_{n=1}^{\infty}$. Then, for small enough $h_{\text{max}}$, there exists a constant $B_{q}>0$ such that
$\left\|\mathcal{F}^{\left(j\right)}\tilde{u}-f^{\left(j\right)}\right\|_{L^{\infty}\left(\Omega^{\left(j\right)}\right)}\leq B_{q}h_{\text{max}}^{q-\frac{1}{2}-\frac{m}{2}}\left(C_{j,q}\left\|u\right\|_{d}+\left\|f^{\left(j\right)}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}\right),$
where $C_{j,q}:=\left\|F^{\left(j,q\right)}d^{-\frac{1}{2}}\right\|_{\ell^{2}}$.
Proof.
We have:
$\left\|\mathcal{F}^{\left(j\right)}\tilde{u}-f^{\left(j\right)}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}$ $\leq\left\|\mathcal{F}^{\left(j\right)}\tilde{u}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}+\left\|f^{\left(j\right)}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}$
$\leq\sum_{n=1}^{\infty}\left\|\mathcal{F}^{\left(j\right)}\phi_{n}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}d_{n}^{-\frac{1}{2}}\left|\tilde{b}_{n}\right|+\left\|f^{\left(j\right)}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}$
[…]
where we note $u$ is a feasible Hermite-Birkhoff interpolant so $\left\|\tilde{b}\right\|_{\ell^{2}}=\left\|\tilde{u}\right\|_{d}\leq\left\|u\right\|_{d}$. By Corollary 7, we then have that there exists some constant $B_{q}$ such that for small enough $h_{\text{max}}$,
$\left\|\mathcal{F}^{\left(j\right)}\tilde{u}-f^{\left(j\right)}\right\|_{L^{\infty}\left(\Omega^{\left(j\right)}\right)}\leq B_{q}h_{\text{max}}^{q-\frac{m}{2}}\left(C_{j,q}\left\|u\right\|_{d}+\left\|f^{\left(j\right)}\right\|_{H^{q}\left(\Omega^{\left(j\right)}\right)}\right)\text{, as required}.$
We note that the $F^{\left(j,q\right)}d^{-\frac{1}{2}}\in\ell^{2}$ condition tells us how quickly our choice of $d^{-\frac{1}{2}}$ must decay. If $\mathcal{F}^{\left(j\right)}$ is order $p$ and has bounded coefficient functions with $q$ bounded derivatives, then $F^{\left(j,q\right)}$ will grow as $\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{p+q}$ as $n\to\infty$.
The error on manifolds of co-dimension greater than zero is slightly more involved since the boundary does not satisfy an interior cone condition. Noting that if $S^{\left(j\right)}$ is a point, then $\mathcal{G}^{\left(j\right)}\tilde{u}=g^{\left(j\right)}$ on that point as long as $S^{\left(j\right)}_{\tilde{M}_{j}}$ is non-empty, we now cover the case that $S^{\left(j\right)}$ has dimension at least one.
Proposition 10.
Assume $S^{\left(j\right)}$ is of dimension $m_{j}>1$ and can be parametrized by a finite atlas $\left\{\sigma^{\left(j\right)}_{k}\right\}_{k=1}^{M}$ where $\sigma^{\left(j\right)}_{k}:U_{k}^{\left(j\right)}\to S^{\left(j\right)}_{k}$ and each $U_{k}^{\left(j\right)}\subset\mathbb{R}^{m_{j}}$ satisfies an interior cone condition. Let $q\in\mathbb{N}$ and assume each $\sigma_{k}^{\left(j\right)}$ is $C^{q}$ with bounded derivatives up to order $q$. Also assume that each $\sigma_{k}^{\left(j\right)}$ has an inverse metric tensor with a bounded norm on $\sigma_{k}^{\left(j\right)}\left(U_{k}^{\left(j\right)}\right)$. Let $j\in\left\{1,2,\ldots,N_{S}\right\}$ be fixed. Also assume $g^{\left(j\right)}$ has bounded derivatives up to order $q$ and $G^{(j,q)}d^{-\frac{1}{2}}\in\ell^{2}$, where $G^{(j,q)}:=\left\{\max_{\left|\beta\right|\leq q}\left\|\partial^{\beta}\mathcal{G}^{\left(j\right)}\phi_{n}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\right\}_{n=1}^{\infty}$. Then, for small enough $h_{\text{max}}$, there exists a constant $\tilde{B}_{q}>0$ such that
$\left\|\mathcal{G}^{\left(j\right)}\tilde{u}-g^{\left(j\right)}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\leq\tilde{B}_{q}h_{\text{max}}^{q-\frac{m_{j}}{2}}\left(\tilde{C}_{j,q}\left\|u\right\|_{d}+\max_{\left|\beta\right|\leq q}\left(\left\|\partial^{\beta}g\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\right)\right),$
where $u\in\mathcal{H}\left(d\right)$ is again a solution to (3.4) and $\tilde{C}_{j,q}:=\left\|G^{\left(j,q\right)}d^{-\frac{1}{2}}\right\|_{\ell^{2}}$.
Proof.
Since the derivatives of $\sigma^{\left(j\right)}_{k}$ are bounded, repeated application of chain rule shows there exist constants $A_{k,q}$ so that for each multi-index $\alpha$ of order less than or equal to $q$,
$\ \left\|\partial^{\alpha}\left(\mathcal{G}^{\left(j\right)}\tilde{u}\circ\sigma^{\left(j\right)}_{k}-g^{\left(j\right)}\circ\sigma^{\left(j\right)}_{k}\right)\right\|_{L^{\infty}\left(U^{(j)}_{k}\right)}$
$\leq$ $\ A_{k,q}\max_{\left|\beta\right|\leq q}\left\|\partial^{\beta}\left(\mathcal{G}^{\left(j\right)}\tilde{u}-g^{\left(j\right)}\right)\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}$
$\leq$ $\ A_{k,q}\max_{\left|\beta\right|\leq q}\left(\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}\tilde{b}_{n}\left\|\partial^{\beta}\mathcal{G}^{\left(j\right)}\phi_{n}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}+\left\|\partial^{\beta}g^{\left(j\right)}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\right)$
[…]
Then there exist constants $\tilde{A}_{k,q}>0$ so that
$\left\|\left(\mathcal{G}^{\left(j\right)}\tilde{u}-g\right)\circ\sigma^{\left(j\right)}_{k}\right\|_{H^{q}\left(U_{k}^{\left(j\right)}\right)}\leq\tilde{A}_{k,q}\left(\tilde{C}_{j,q}\left\|\tilde{b}\right\|_{\ell^{2}}+\max_{\left|\beta\right|\leq q}\left(\left\|\partial^{\beta}g^{\left(j\right)}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\right)\right).$
Once again using Corollary 7, we have that there exist constants $\tilde{B}_{k,q}>0$ so that, for small enough $h_{\text{max}}$,
$\left\|\left(\mathcal{G}^{\left(j\right)}\tilde{u}-g\right)\circ\sigma^{\left(j\right)}_{k}\right\|_{L^{\infty}\left(U_{k}^{\left(j\right)}\right)}\leq\tilde{B}_{k,q}h_{\text{max}}^{q-\frac{m_{j}}{2}}\left(\tilde{C}_{j,q}\left\|u\right\|_{d}+\max_{\left|\beta\right|\leq q}\left(\left\|\partial^{\beta}g^{\left(j\right)}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\right)\right).$
Note that Corollary 7 is applied to $U^{\left(j\right)}_{k}$, not $S^{\left(j\right)}_{k}$, so we are implicitly relying on the boundedness of the inverse metric tensor of $\sigma^{\left(j\right)}_{k}$ so that the fill distance on $U^{\left(j\right)}_{k}$ is at most a constant multiple of the fill distance on $S^{\left(j\right)}_{k}$.
Finally, letting $\tilde{B}_{q}=\max\left\{\tilde{B}_{k,q}\right\}_{k=1}^{M}$, we have the desired result:
$\left\|\mathcal{G}^{\left(j\right)}\tilde{u}-g^{\left(j\right)}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\leq\tilde{B}_{q}h_{\text{max}}^{q-\frac{m_{j}}{2}}\left(\tilde{C}_{j,q}\left\|u\right\|_{d}+\max_{\left|\beta\right|\leq q}\left(\left\|\partial^{\beta}g^{\left(j\right)}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\right)\right).$
3.6. Finite Number of Basis Functions
We now truncate our Fourier series to a finite basis of $N_{b}$ terms. Let $\mathcal{L}_{N_{b}}:=\mathcal{L}_{\text{HB}}\biggr{|}_{\mathbb{C}^{N_{b}}}:\mathbb{C}^{N_{b}}\to\mathbb{C}^{\tilde{N}}$, where $\mathcal{L}_{\text{HB}}$ is as defined in the proof of Proposition 8. Let $\boldsymbol{f}:=\left(f^{\left(j\right)}\left(\boldsymbol{x}_{k}^{\left(j\right)}\right)\right)$, $\boldsymbol{g}:=\left(g^{\left(j\right)}\left(\boldsymbol{y}_{k}^{\left(j\right)}\right)\right)$ be vectorized forms of all Hermite-Birkhoff data from (3.4) (column vectors using all values of $j$ and $k$). The problem then becomes:
[…]
(3.9) subject to: $\mathcal{L}_{N_{b}}\boldsymbol{b}=\text{proj}_{\mathcal{R}\left(\mathcal{L}_{N_{b}}\right)}\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right).$
If a Hermite-Birkhoff interpolant exists in $\text{span}\left\{\phi_{n}\right\}_{n=1}^{N_{b}}$, then
$\text{proj}_{\mathcal{R}\left(\mathcal{L}_{N_{b}}\right)}\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right)=\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right).$
Let $\tilde{\boldsymbol{b}}$ be the solution to this optimization problem.
Proposition 11.
Assume there is a constant $B_{p}$ such that for all large enough $n$, $\left\|\mathcal{F}^{\left(j\right)}\phi_{n}\right\|_{L^{\infty}\left(\Omega^{\left(j\right)}\right)}\leq B_{p}\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{p}$ for each $j\in\left\{1,2,\ldots,N_{\Omega}\right\}$ and $\left\|\mathcal{G}^{\left(j\right)}\phi_{n}\right\|_{L^{\infty}\left(S^{\left(j\right)}\right)}\leq B_{p}\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{p}$ for each $j\in\left\{1,2,\ldots,N_{S}\right\}$. Also assume $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$. Then, for a fixed number of total Hermite-Birkhoff interpolation conditions $\tilde{N}$, there exists a constant $A_{\tilde{N}}>0$ such that for sufficiently large $N_{b}$,
$\left\|\tilde{u}_{\infty}-\tilde{u}_{N_{b}}\right\|_{L^{\infty}\left(\Omega\right)}\leq A_{\tilde{N}}\left(\sum_{n=N_{b}+1}^{\infty}\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{2p}d_{n}^{-1}\right)^{\frac{1}{2}},$
where $\tilde{u}_{\infty}:=\sum_{n=1}^{N_{b}}d_{n}^{-\frac{1}{2}}\tilde{b}_{\infty,n}\phi_{n}\in\mathcal{H}\left(d\right)$ is from the solution $\tilde{b}_{\infty}\in\ell^{2}$ to the full $\ell^{2}$ problem (3.5) and $\tilde{u}_{N_{b}}$ is the finite basis solution: $\tilde{u}_{N_{b}}=\sum_{n=1}^{N_{b}}d_{n}^{-\frac{1}{2}}\tilde{b}_{n}\phi_{n}$, where $\tilde{b}_{n}$ is from the solution $\tilde{\boldsymbol{b}}$ to (3.6).
The approach of the proof is to note that if an interpolant exists for finite $N_{b}$, it is also feasible for the $\ell^{2}$ problem, while the truncated solution to the $\ell^{2}$ problem will “nearly” be an interpolant and will be close to a feasible solution for the finite basis problem. Along with orthogonality of the minimizer to the null space, this turns out to be enough to show convergence proportional to $\left(\sum_{n=N_{b}+1}^{\infty}\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{2p}d_{n}^{-1}\right)^{\frac{1}{2}}$ of the truncated basis solution to the $\ell^{2}$ problem. The idea here is to show that adding terms to our Fourier basis recovers the solution to the full $\ell^{2}$ problem in the limit $N_{b}\to\infty$.
Proof.
To start, let $\tilde{N}$ be fixed. Let $\boldsymbol{V}_{N_{b}}$ be the matrix corresponding to $\mathcal{L}_{N_{b}}$, so that
$\boldsymbol{V}_{N_{b}}\tilde{\boldsymbol{b}}$ $=\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right),$
then the unique solution to the finite-dimensional problem is
$\tilde{\boldsymbol{b}}=\boldsymbol{V}_{N_{b}}^{+}\left(\begin{array}[]{cc}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right),$
where + indicates the Moore-Penrose pseudoinverse. Now, let $N_{b,\text{interp}}$ be the smallest number of basis functions for which $\mathcal{L}_{N_{b}}$ is surjective (such a finite value always exists, see Remark 2). Note that if $\tilde{\boldsymbol{b}}_{N_{b,\text{interp}}}$ is the solution to the optimization problem for $N_{b,\text{interp}}$ basis functions, then $\tilde{\boldsymbol{b}}_{N_{b,\text{interp}}}$ padded with zeros to be in $\mathbb{C}^{N_{b}}$ is feasible for any $N_{b}>N_{b,\text{interp}}$. Therefore,
$\left\|\tilde{\boldsymbol{b}}\right\|_{2}\leq\left\|\tilde{\boldsymbol{b}}_{N_{b,\text{interp}}}\right\|_{2}\text{, for any }N_{b}>N_{b,\text{interp}}.$
This holds for any $\boldsymbol{f},\boldsymbol{g}$, so if we define $M_{\tilde{N}}=\left\|\boldsymbol{V}_{N_{b,\text{interp}}}^{+}\right\|_{2}$ , then as long as $N_{b}\geq N_{b,\text{interp}}$,
$\left\|\boldsymbol{V}_{N_{b}}^{+}\right\|_{2}\leq M_{\tilde{N}}.$
In particular, $\left\|\boldsymbol{V}_{N_{b}}^{+}\right\|_{2}$ is bounded as $N_{b}\to\infty$. Now define $\tilde{b}_{\infty}$ as the solution to the full $\ell^{2}$ problem, and define the truncated $\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}=\left(\tilde{b}_{\infty,n}\right)_{n=1}^{N_{b}}$. Then, define
$\varepsilon_{\text{trunc}}:=\left\|\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right)-\boldsymbol{V}_{N_{b}}\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}\right\|_{2},$
and note that
$\boldsymbol{V}_{N_{b}}\left(\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}+\boldsymbol{V}_{N_{b}}^{+}\left(\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right)-\boldsymbol{V}_{N_{b}}\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}\right)\right)=\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right).$
Therefore, since $\tilde{\boldsymbol{b}}$ is the minimizer,
(3.12) $\left\|\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}+\boldsymbol{V}_{N_{b}}^{+}\left(\left(\begin{array}[]{c}\boldsymbol{f}\\ \boldsymbol{g}\end{array}\right)-\boldsymbol{V}_{N_{b}}\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}\right)\right\|_{2}$ $\geq\left\|\tilde{\boldsymbol{b}}\right\|_{2}$
(3.13) $\implies M_{\tilde{N}}\varepsilon_{\text{trunc}}\geq\left\|\tilde{\boldsymbol{b}}\right\|_{2}-\left\|\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}\right\|_{2}$ $\geq\left\|\tilde{\boldsymbol{b}}\right\|_{2}-\left\|\tilde{b}_{\infty}\right\|_{\ell^{2}}\geq 0,$
where we note $\left\|\tilde{\boldsymbol{b}}\right\|_{2}\geq\left\|\tilde{b}_{\infty}\right\|_{\ell^{2}}\geq\left\|\tilde{\boldsymbol{b}}_{\infty}^{\left(N_{b}\right)}\right\|_{2}$ since $\tilde{\boldsymbol{b}}$ padded with zeros is feasible for the full $\ell^{2}$ problem. More precisely, let $\mathcal{E}:\mathbb{C}^{N_{b}}\to\ell^{2}$ so that $\left(\mathcal{E}\boldsymbol{b}\right)_{n}=b_{n}$ for $n\leq N_{b}$ and zero otherwise, then $\mathcal{E}\tilde{\boldsymbol{b}}$ is feasible for the $\ell^{2}$ problem and $\left\|\tilde{\boldsymbol{b}}\right\|_{2}=\left\|\mathcal{E}\tilde{b}\right\|_{\ell^{2}}\geq\left\|\tilde{b}_{\infty}\right\|_{\ell^{2}}$. $\tilde{b}_{\infty}$ is the minimizer and is therefore orthogonal to the null space of $\mathcal{L}_{\text{HB}}$, so $\left\|\tilde{b}_{\infty}\right\|_{\ell^{2}}^{2}=\left(\tilde{b}_{\infty},\mathcal{E}\tilde{\boldsymbol{b}}\right)_{\ell^{2}}$ and
$\left\|\tilde{b}_{\infty}-\mathcal{E}\tilde{\boldsymbol{b}}\right\|_{\ell^{2}}^{2}=\left\|\tilde{\boldsymbol{b}}\right\|_{2}^{2}-\left\|\tilde{b}_{\infty}\right\|_{\ell^{2}}^{2}\leq\left(\left\|\tilde{b}_{\infty}\right\|_{\ell^{2}}+\left\|\tilde{\boldsymbol{b}}\right\|_{2}\right)M_{\tilde{N}}\varepsilon_{\text{trunc}},$
where we use equation (3.13). Then,
$\left\|\tilde{b}_{\infty}\right\|_{\ell^{2}}+\left\|\tilde{\boldsymbol{b}}\right\|_{2}$ $=\left\|\tilde{b}_{\infty}\right\|_{\ell^{2}}+\left\|\tilde{\boldsymbol{b}}\right\|_{2}-\left\|\tilde{b}_{\infty}\right\|_{\ell^{2}}+\left\|\tilde{b}_{\infty}\right\|_{\ell^{2}}$
$\leq 2\left\|\tilde{b}_{\infty}\right\|_{\ell^{2}}+M_{\tilde{N}}\varepsilon_{\text{trunc}}$
$\implies\left\|\tilde{b}_{\infty}-\mathcal{E}\tilde{\boldsymbol{b}}\right\|_{\ell^{2}}^{2}$ $\leq\left(2\left\|\tilde{b}_{\infty}\right\|_{\ell^{2}}+M_{\tilde{N}}\varepsilon_{\text{trunc}}\right)M_{\tilde{N}}\varepsilon_{\text{trunc}},$
where we again use (3.13). Now, using the line above and Cauchy-Schwarz,
$\left\|\tilde{u}_{\infty}-\tilde{u}_{N_{b}}\right\|_{L^{\infty}\left(\Omega\right)}$ $=\left\|\left(\sum_{n=1}^{\infty}d_{n}^{-\frac{1}{2}}\tilde{b}_{\infty,n}\phi_{n}\right)-\left(\sum_{n=1}^{N_{b}}d_{n}^{-\frac{1}{2}}\tilde{b}_{n}\phi_{n}\right)\right\|_{L^{\infty}\left(\Omega\right)}$
$\leq\left(\left|\tilde{b}_{\infty}-\mathcal{E}\tilde{\boldsymbol{b}}\right|,d_{n}^{-\frac{1}{2}}\right)_{\ell^{2}}$
(3.14) $\leq\left\|d^{-\frac{1}{2}}\right\|_{\ell^{2}}\sqrt{\left(2\left\|\tilde{b}_{\infty}\right\|_{\ell^{2}}+M_{\tilde{N}}\varepsilon_{\text{trunc}}\right)M_{\tilde{N}}\varepsilon_{\text{trunc}}}.$
Also,
$\varepsilon_{\text{trunc}}^{2}$ $=\sum_{j=1}^{N_{\Omega}}\sum_{k=1}^{\tilde{N}_{j}}\left|f^{\left(j\right)}\left(\boldsymbol{x}^{\left(j\right)}_{k}\right)-\sum_{n=1}^{N_{b}}d_{n}^{-\frac{1}{2}}\tilde{b}_{\infty,n}\mathcal{F}^{\left(j\right)}\phi_{n}\left(\boldsymbol{x}^{\left(j\right)}_{k}\right)\right|^{2}$
$\quad+\sum_{j=1}^{N_{S}}\sum_{k=1}^{\tilde{M}_{j}}\left|g^{\left(j\right)}\left(\boldsymbol{y}^{\left(j\right)}_{k}\right)-\sum_{n=1}^{N_{b}}d_{n}^{-\frac{1}{2}}\tilde{b}_{\infty,n}\mathcal{G}^{\left(j\right)}\phi_{n}\left(\boldsymbol{y}^{\left(j\right)}_{k}\right)\right|^{2}$
$=\sum_{j=1}^{N_{\Omega}}\sum_{k=1}^{\tilde{N}_{j}}\left|\sum_{n=N_{b}+1}^{\infty}d_{n}^{-\frac{1}{2}}\tilde{b}_{\infty,n}\mathcal{F}^{\left(j\right)}\phi_{n}\left(\boldsymbol{x}^{\left(j\right)}_{k}\right)\right|^{2}$
$\quad+\sum_{j=1}^{N_{S}}\sum_{k=1}^{\tilde{M}_{j}}\left|\sum_{n=N_{b}+1}^{\infty}d_{n}^{-\frac{1}{2}}\tilde{b}_{\infty,n}\mathcal{G}^{\left(j\right)}\phi_{n}\left(\boldsymbol{y}^{\left(j\right)}_{k}\right)\right|^{2}$
[…]
where the last line holds for large enough $N_{b}$ by assumption on $\mathcal{F}^{\left(j\right)},\mathcal{G}^{\left(j\right)}$. Finally, from (3.14),
$\left\|\tilde{u}_{\infty}-\tilde{u}_{N_{b}}\right\|_{L^{\infty}\left(\Omega\right)}$
$\leq$ $\left\|d^{-\frac{1}{2}}\right\|_{\ell^{2}}\left(2\left\|\tilde{b}_{\infty}\right\|_{\ell^{2}}+M_{\tilde{N}}B_{p}\sqrt{\tilde{N}}\left(\sum_{n=N_{b}+1}^{\infty}\left|\tilde{b}_{\infty,n}\right|^{2}\right)^{\frac{1}{2}}\left(\sum_{n=N_{b}+1}^{\infty}\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{2p}d_{n}^{-1}\right)^{\frac{1}{2}}\right)^{\frac{1}{2}}$
$\quad\cdot\left(M_{\tilde{N}}B_{p}\sqrt{\tilde{N}}\left(\sum_{n=N_{b}+1}^{\infty}\left|\tilde{b}_{\infty,n}\right|^{2}\right)^{\frac{1}{2}}\left(\sum_{n=N_{b}+1}^{\infty}\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{2p}d_{n}^{-1}\right)^{\frac{1}{2}}\right)^{\frac{1}{2}}.$
Now, from equation (3.3) in Subsection 3.3, we recall $\tilde{b}_{\infty}\in\mathcal{R}\left(\mathcal{L}_{\text{HB}}^{*}\right)$. This means $\left|\tilde{b}_{\infty,n}\right|=\mathcal{O}\left(\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{p}d_{n}^{-\frac{1}{2}}\right)$ as $n\to\infty$. From the bound above, we can conclude that for large enough $N_{b}$, there exists some constant $A_{\tilde{N}}$ such that
$\left\|\tilde{u}_{\infty}-\tilde{u}_{N_{b}}\right\|_{L^{\infty}\left(\Omega\right)}\leq A_{\tilde{N}}\left(\sum_{n=N_{b}+1}^{\infty}\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{2p}d_{n}^{-1}\right)^{\frac{1}{2}}.$
That is, $\tilde{u}_{N_{b}}\to\tilde{u}_{\infty}$ uniformly on $\Omega$ as long as $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$. ∎
The $A_{\tilde{N}}$ coefficient of the Proposition 11 could be problematic in theory, but in practice, we can make $\left(\sum_{n=N_{b}+1}^{\infty}\left\|\boldsymbol{\omega}_{n}\right\|_{2}^{2p}d_{n}^{-1}\right)^{\frac{1}{2}}$ converge arbitrarily quickly in $N_{b}$ with a suitable choice of $d_{n}$, so any highly underdetermined system will produce reasonable results. Furthermore, $M_{\tilde{N}}$, which is a term in $A_{\tilde{N}}$, could be replaced by $\left\|\boldsymbol{V}_{N_{b}}^{+}\right\|_{2}$, which is non-increasing and will typically decrease as $N_{b}$ increases.
Remark 2.
In 1D, an interpolant for a combination of function or derivative interpolation conditions (a Hermite-Birkhoff interpolant) exists under certain conditions when $N_{b}=\tilde{N}$ [johns75], but this is not necessarily the case in dimensions greater than one, again due to the Mairhuber-Curtis Theorem [curti59, mairh56]. An interpolant does, however, always exist with some finite number of basis functions.
In particular, it can be quickly shown that $N_{b,\text{interp}}\leq\left(\left(p+1\right)\tilde{N}_{\text{distinct}}+1\right)^{m}$, where $\tilde{N}_{\text{distinct}}\leq\tilde{N}$ is the number of distinct points in the box $\Omega\subset\mathbb{R}^{m}$. $\tilde{N}_{\text{distinct}}$ can be less than $\tilde{N}$ if one or more points have two or more interpolation conditions imposed. For $m=1$, it has been known for some time [johns75, Thm. 1] that we can always construct a trigonometric polynomial with at most $\left(p+1\right)\tilde{N}_{\text{distinct}}+1$ terms that is zero and has all derivatives of up to order $p$ equal to zero on up to $\tilde{N}_{\text{distinct}}$ distinct points of our choosing, except for one chosen point where either the function or one of its derivatives are equal to one. For $m>1$, we can take products of these 1D functions to get a trigonometric polynomial with at most $\left(\left(p+1\right)\tilde{N}_{\text{distinct}}+1\right)^{m}$ terms that is zero and has derivatives equal to zero on a tensor grid of $\tilde{N}_{\text{distinct}}^{m}$ points (that our desired $\tilde{N}_{\text{distinct}}$ points are a subset of), except for one point and derivative of interest. A linear combination of these functions is a feasible interpolant. Experimentally, however, we typically observe $N_{b,\text{interp}}$ to simply be the total number of interpolation conditions ($\tilde{N}$) or not much larger (less than $2\tilde{N}$).
Theorem 2.2 of [chand18] provides an upper bound on $N_{b,\text{interp}}$ that depends on point separation and will often be less than the simple bound presented above. Note that increasing $N_{b}$ well beyond $N_{b,\text{interp}}$ can improve the error, in accordance with Proposition 11.
3.7. Boundedness, Solvability, and Convergence
In many applications, we are not only interested in finding solutions to a PDE; we may also want to first determine the values of parameters for which the PDE has a solution at all. These are existence problems, which can often be difficult to analyze numerically. Of particular interest in numerical analysis are eigenvalue problems, where we can find the eigenvalues of $\mathcal{F}$ by finding the values of $\lambda\in\mathbb{C}$ for which $\left(\mathcal{F}-\lambda\right)u=0$ has a non-zero solution.
Consider, for some linear differential operators $\mathcal{F,G}$ of order at most $p$, the PDE:
(3.15) $\mathcal{F}u=f\text{, on }S,\quad\mathcal{G}u\biggr{|}_{\partial S}=g.$
To discretize this problem, let $S_{\tilde{N}}:=\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\tilde{N}}\subset S$ and $\partial S_{\tilde{N}_{\partial}}:=\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\tilde{N}_{\partial}}\subset\partial S$ be sets of points. Then, let $\boldsymbol{a}\in S$ be fixed. The discretized problem is then
(3.16) minimize, over $u^{\left(\tilde{N}\right)}\in\mathcal{H}\left(d\right)$: $\left\|u^{\left(\tilde{N}\right)}\right\|_{d}$
subject to: $\mathcal{F}u^{\left(\tilde{N}\right)}\biggr{|}_{S_{\tilde{N}}}=f,\ \ \mathcal{G}u^{\left(\tilde{N}\right)}\biggr{|}_{\partial S_{\tilde{N}_{\partial}}}=g,\ u^{\left(\tilde{N}\right)}\left(\boldsymbol{a}\right)=1,$
where we only include the $u^{\left(\tilde{N}\right)}\left(\boldsymbol{a}\right)=1$ condition in the case $f=g=0$ to ensure we get a non-zero solution. If there is a non-zero solution $u$ to the original PDE in $\mathcal{H}\left(d\right)$ (possibly requiring $u\left(\boldsymbol{a}\right)=1$ as well), then $u$ is feasible, and $\left\|u^{\left(\tilde{N}\right)}\right\|_{d}\leq\left\|u\right\|_{d}$ for all $\tilde{N}$. We have shown convergence of $\mathcal{F}u^{\left(\tilde{N}\right)}\to\mathcal{F}u$ in this case (Proposition 9). However, particularly for eigenvalue problems, we may also be interested in sufficient conditions for a solution to exist in $\mathcal{H}\left(d\right)$, and for $u^{\left(\tilde{N}\right)}$ to converge to a solution.
In Theorem 7.1 of the work by Chandrasekaran, Gorman, and Mhaska [chand18], the authors prove a result that, in our notation, would show $\mathcal{F}u\to f$ and $\mathcal{G}u\to g$ pointwise when a solution to (3.15) exists. Our next result adds to this for our approach by noting that mere boundedness of $u^{\left(\tilde{N}\right)}$ (which would be implied by feasibility) is enough to imply that a solution to (3.15) must exist, and that $u^{\left(\tilde{N}\right)}$ will converge to a solution $u$ in the $d$-norm. For choices of $d$ such that $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$, this implies uniform convergence of $u^{\left(\tilde{N}\right)}$ and its derivatives of up to order $p$ to $u$ over all of $\Omega$.
Proposition 12.
Let $\mathcal{F,G}$ be $p$-th or lower order linear differential operators with bounded coefficient functions, and assume $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$. Let $S_{1}\subset S_{2}\subset\cdots\subset S$, where $S_{n}=\left\{\boldsymbol{x}_{j}\right\}_{j=1}^{\tilde{N}_{n}}$ such that $\left\{\tilde{N}_{n}\right\}$ is strictly increasing. Let $\partial S_{1}\subset\partial S_{2}\subset\cdots\subset\partial S$, where $\partial S_{n}=\left\{\boldsymbol{y}_{j}\right\}_{j=1}^{\tilde{N}_{\partial,n}}$ such that $\left\{\tilde{N}_{\partial,n}\right\}$ is strictly increasing. Let $u^{\left(n\right)}$ be the corresponding solution to (3.16) on $S_{n}$ and $\partial S_{n}$. If, for all $n$, $\left\|u^{\left(n\right)}\right\|_{d}\leq\tilde{Q}$ for some $\tilde{Q}>0$, and $h_{\text{max}}^{\left(n\right)}\to 0$, where $h_{\text{max}}^{\left(n\right)}$ is the fill distance associated with $S_{n}$ and $\partial S_{n}$, then there exists $u\in\mathcal{H}\left(d\right)$ so that $u\biggr{|}_{S}$ is a solution to (3.15) and $u^{\left(n\right)}\to u$ in $\mathcal{H}\left(d\right)$ and all derivatives of $u^{\left(n\right)}$ up to order $p$ converge uniformly to the derivatives of $u$ on $\Omega$.
Proof.
First, note that for all $m>n$, $u^{\left(m\right)}$ is feasible for (3.16) on $S_{n},\partial S_{n}$, since $S_{n}\subset S_{m},\partial S_{n}\subset\partial S_{m}$. The fact that $u^{\left(n\right)}$ is the solution with minimum $d$-norm implies $\left\{\left\|u^{\left(n\right)}\right\|\right\}_{n}$ is a non-decreasing sequence. $\left\{\left\|u^{\left(n\right)}\right\|\right\}_{n}$ is bounded above by $\tilde{Q}$, and therefore $\left\|u^{\left(n\right)}\right\|_{d}\uparrow Q$ for some $Q>0$ ($u^{\left(n\right)}$ cannot be zero since either $u^{\left(n\right)}\left(\boldsymbol{a}\right)=1$ or one of $f,g$ is non-zero, and $u^{\left(n\right)}$ is $C^{p}$ since $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$ implies uniform convergence of the $u^{\left(n\right)}$ Fourier series and its derivatives of up to order $p$).
Let $\varepsilon>0$. Then, since, $\left\|u^{\left(n\right)}\right\|_{d}\uparrow Q$, there exists some $M>0$ so that for all $n,m>M$ such that $n<m$,
(3.17) $0\leq\left\|u^{\left(m\right)}\right\|_{d}-\left\|u^{\left(n\right)}\right\|_{d}<\frac{\varepsilon^{2}}{2Q}.$
Now, recall that $u^{\left(m\right)}$ is feasible for (3.16) on $S_{n},\partial S_{n}$, since $S_{n}\subset S_{m}$ and $\partial S_{n}\subset\partial S_{m}$, so
$\left(u^{\left(m\right)}-u^{\left(n\right)},u^{\left(n\right)}\right)_{d}$ $=0\text{, since $u^{\left(n\right)}$is the minimal solution}$
(3.18) $\implies\left(u^{\left(m\right)},u^{\left(n\right)}\right)_{d}=\left(u^{\left(n\right)},u^{\left(m\right)}\right)_{d}$ $=\left\|u^{\left(n\right)}\right\|_{d}^{2}.$
Using (3.18) and then (3.17), it can then be quickly shown that
$\left\|u^{\left(n\right)}-u^{\left(m\right)}\right\|_{d}$ $\leq\varepsilon.$
Therefore, $\left\{u^{\left(n\right)}\right\}$ is Cauchy and must converge to some $u$ in the $d$-norm since $\mathcal{H}\left(d\right)$ is a Hilbert space. Let $\hat{u}^{\left(n\right)}$ and $\hat{u}$ be the Fourier coefficients of $u^{\left(n\right)}$ and $u$, respectively. Then, for all $q\leq p$, and for any $q$-th order multi-index $\alpha$,
$\left\|\partial^{\alpha}\left(u^{\left(n\right)}-u\right)\right\|_{L^{\infty}\left(\Omega\right)}$ $\leq\sum_{j=1}^{\infty}\left\|\boldsymbol{\omega}_{j}\right\|_{2}^{q}d^{-\frac{1}{2}}d^{\frac{1}{2}}\left|\hat{u}_{j}^{\left(n\right)}-\hat{u}_{j}\right|$
[…]
$\to 0\text{, as $n\to\infty$}.$
That is, partial derivatives of $u^{\left(n\right)}$ up to order $p$ converge uniformly to the partial derivatives of $u$. This implies $\mathcal{F}u=f$ on $\bigcup_{n=1}^{\infty}S_{n}$. Then, since $h_{\text{max}}^{\left(n\right)}\to 0$, $\bigcup_{n=1}^{\infty}S_{n}$ is dense in $S$, and by continuity of the partial derivatives of $u$ up to order $p$ (again implied by $\left\|\boldsymbol{\omega}\right\|_{2}^{p}d^{-\frac{1}{2}}\in\ell^{2}$), $\mathcal{F}u=f$ on $S$. A similar argument can be applied on the boundary to show $\mathcal{G}u\biggr{|}_{\partial S}=g$. Therefore, $u$ solves (3.15) and $u^{\left(n\right)}\to u$ in the $d$-norm (and uniformly in all derivatives up to order $p$). ∎
Note that we could impose additional conditions on additional domains and reach a similar result. Also notice that if $\mathcal{F}=\mathcal{K}-\lambda$ and $g=0$, then $u$ is an eigenfunction of $\mathcal{K}$ with eigenvalue $\lambda$, subject to the given boundary condition.
A particular application of Proposition 12 is in investigating solvability. Informally, Proposition 12 shows:
[…]
and we already know:
[…]
We can then conclude that there is an equivalence:
[…]
This means that for a dense enough point cloud, the $d$-norm of interpolant solutions to solvable PDEs will be much smaller than the $d$-norm for solutions for PDEs without a solution; if the PDE is not solvable, the norm must be unbounded. We test this numerically in Subsection 4.2 to search for eigenvalues.
4. Numerical Experiments
4.1. Surface Poisson
We now move on to testing convergence for a surface PDE. Specifically, we solve a Poisson problem on a catenoid with a “wavy” edge. The surface is given by:
$S=\left\{\left(\cosh\left(t\right)\cos\left(s\right),\cosh\left(t\right)\sin\left(s\right),t\right):s\in\left[0,2\pi\right),\left(t-0.1\sin\left(3s\right)\right)\in\left[-1,1\right]\right\}.$
We solve the Poisson problem:
$-\Delta_{S}u\left(s,t\right)$ $=16\frac{\cos\left(4s\right)}{\cosh^{2}\left(t\right)}=:f\left(s,t\right),\ u\left(s,t\right)\biggr{|}_{\partial S}=\cos\left(4s\right)=:g\left(s,t\right).$
The exact solution to this problem is $u\left(s,t\right)=\cos\left(4s\right)$. To solve this, we impose Hermite-Birkhoff interpolation conditions on a point cloud $S_{\tilde{N}}$ and boundary point cloud $\partial S_{\tilde{N}_{\partial}}$ to attempt to find a first-order Closest Point-like extension that approximately solves the PDE (see Proposition 3). Note that such an extension exists due to Proposition 5. Our discretized problem is:
(4.1) minimize: $\left\|\boldsymbol{b}\right\|_{2}$
subject to: $\sum_{n=1}^{N_{b}}d_{n}^{-\frac{1}{2}}b_{n}\left(\Delta\phi_{n}-\hat{\boldsymbol{n}}^{*}\left(D^{2}\phi_{n}\right)\hat{\boldsymbol{n}}\right)=f\text{, on $S_{\tilde{N}}$}$
$\sum_{n=1}^{N_{b}}d_{n}^{-\frac{1}{2}}b_{n}\hat{\boldsymbol{n}}^{*}\nabla\phi_{n}=0\text{, on $S_{\tilde{N}}$},\ \sum_{n=1}^{N_{b}}d_{n}^{-\frac{1}{2}}b_{n}\phi_{n}=g\text{, on $\partial S_{\tilde{N}_{\partial}}$}.$
Our choice of $d_{n}$ is $\left(\exp\left(q\sqrt{2\pi/T}\right)+\exp\left(q\sqrt{\left\|\boldsymbol{\omega}_{n}\right\|_{2}}\right)\right)^{2}$ with $q=4$; this choice achieves super-algebraic convergence when there is a solution in the native space. This is since the sequences $F^{\left(j,q\right)}d^{-\frac{1}{2}}$ and $G^{\left(j,q\right)}d^{-\frac{1}{2}}$ from Propositions 9 and 10, respectively, will both be in $\ell^{2}$ for all $q$ for this problem and choice of $d$. $T$ can be used to control the oscillation width of the $\psi$ functions from Subsection 3.3. We use $T=2$, which will provide fast convergence at the cost of poorer conditioning for dense point clouds, compared to smaller values of $T$. $T$ serves the same purpose as RBF shape parameters; however, conditioning here is not as much of an issue compared to most RBF methods, meaning a wider range of parameters are suitable. We discuss this further at the end of this subsection.
The other parameters used are $\ell=4$ for a $4\times 4\times 4$ extension domain $\Omega$, $N_{b}=\left(2\cdot 13+1\right)^{3}$ Fourier basis functions, and $\tilde{N}_{s}$ unique $s$ values so that $h_{\text{max}}\propto\tilde{N}_{s}^{-1}$. A plot of the computed solution for $\tilde{N}_{s}=100$ is in Figure 1 and a convergence test recorded in Table 1.
Figure 1. A plot of the surface Poisson solution for $\tilde{N}_{s}=100$.
$\tilde{N}_{s}$ Max Error Convergence Rate
20 $3.4619\times 10^{-3}$ N/A
40 $4.6741\times 10^{-5}$ 6.211
60 $1.7315\times 10^{-6}$ 8.128
80 $5.1884\times 10^{-8}$ 12.193
100 $3.1698\times 10^{-9}$ 12.527
Table 1. Recorded max error and estimated convergence rate on the point cloud for the surface Poisson problem for various $\tilde{N}_{s}$ values.
We see high-order convergence in Table 1. An important note is that such a low error for $\tilde{N}_{s}=100$ is not possible without solving the optimization problem directly; we use a complete orthogonal decomposition. Forming the kernel matrix $\boldsymbol{\Phi}_{N_{b}}=\boldsymbol{V}_{N_{b}}\boldsymbol{V}_{N_{b}}^{*}$ results in a much more poorly conditioned linear system to solve, akin to direct RBF methods, causing convergence to stall. Note that the condition number of $\boldsymbol{V}_{N_{b}}\boldsymbol{V}_{N_{b}}^{*}$ is the square of the condition number of $\boldsymbol{V}_{N_{b}}$. The error for $\tilde{N}_{s}=100$ that is obtained when solving the system with the kernel matrix naively is much larger at $4.7372\times 10^{-6}$. Having additional, more stable options for finding a solution is a benefit of underdetermined Fourier extensions over Hermite RBFs, while the option to form the kernel matrix for the Fourier extensions also remains.
4.2. Eigenvalue Problem
As mentioned in Subsection 3.7, we may be able to use the fact that the $d$-norm of interpolant solutions is bounded if and only if a solution exists to investigate solvability numerically. As an example, we can consider the eigenvalue problem on the unit sphere $S=\mathbb{S}^{2}$ for the (negative) Laplace-Beltrami operator:
$-\Delta_{S}u-\lambda u=0\text{, for some $u\neq 0$}.$
The eigenvalues $\lambda$ are $n\left(n+1\right)$ for non-negative integers $n$. To find the eigenvalues numerically, we can solve the problem:
(4.2) $\left(-\Delta\tilde{u}-\hat{\boldsymbol{n}}_{S}\cdot\left(D^{2}\tilde{u}\right)\hat{\boldsymbol{n}}_{S}-\lambda\right)\biggr{|}_{S_{\tilde{N}}}=0,\quad\hat{\boldsymbol{n}}_{S}\cdot\nabla\tilde{u}\biggr{|}_{S_{\tilde{N}}}=0,\quad\tilde{u}\left(0,0,1\right)=1.$
where $S_{\tilde{N}}$ is a set of $\tilde{N}$ randomly placed points on the sphere, and we choose the interpolant $\tilde{u}=\sum_{n=1}^{N_{b}}d_{n}^{-\frac{1}{2}}\tilde{b}_{n}\phi_{n}$ such that $\left\|\tilde{\boldsymbol{b}}\right\|_{2}$ is minimized subject to the constraints in (4.2).
The final condition ensures we produce a non-zero solution. At first glance, the location of the single point $\boldsymbol{a}$ where we enforce $\tilde{u}\left(\boldsymbol{a}\right)=1$ may seem to be important, since the eigenspace for a single eigenvalue may have $\tilde{u}\left(\boldsymbol{a}\right)=0$ for our selected $\boldsymbol{a}$. However, a random point would work with probability one, since the zero set of any eigenfunction is measure zero. That is, for arbitrary surfaces, choosing a point randomly should not cause an issue in practice. In this sphere example, symmetry ensures that there is an eigenfunction with $\tilde{u}\left(0,0,1\right)=1$ for any eigenvalue of $-\Delta_{S}$, and any point location would work. Another condition that forces the function to be non-zero would also be acceptable, such as setting a sum of function values or derivatives at random points to be equal to 1, but in our testing, setting $\tilde{u}\left(\boldsymbol{a}\right)=1$ for a single point $\boldsymbol{a}$ is sufficient. Eigenvalue multiplicities can be explored by setting multiple points where the function must be non-zero, but we leave this for later work.
The $d$-norm of the interpolant for various values of $\lambda$ is shown in Figure 2, along with red lines at the true eigenvalues.
Figure 2. $d$-norm of interpolant solution for various $\lambda$ ($\tilde{N}=400$ randomly generated points on the sphere, $25^{3}$ Fourier basis functions on $\Omega=\left(-2,2\right)^{3}$, $d_{n}=\left(\exp\left(4\sqrt{2\pi/T}\right)+\exp\left(4\sqrt{\left\|\boldsymbol{\omega}_{n}\right\|_{2}}\right)\right)^{2}$, $T=4$). True eigenvalues are indicated by red lines.
We see that the minima of the $d$-norm align with the correct eigenvalues in Figure 2. We also look specifically at the convergence of the first non-zero eigenvalue ($\lambda=2$) in Table 2; we use a simple bisection-like search to find the local minimum. $h_{\text{avg}}$ is the average distance between points on the surface and their closest points in the point cloud and is inversely proportional to $\sqrt{\tilde{N}}$.
$\tilde{N}$ $\left|\lambda_{\text{est}}-2\right|$ Convergence Rate ($\tilde{N}$) Convergence Rate ($h_{\text{avg}}$)
100 $1.7868\times 10^{-2}$ N/A N/A
150 $3.3266\times 10^{-3}$ 4.146 8.292
200 $1.1514\times 10^{-4}$ 11.692 23.384
250 $9.2243\times 10^{-6}$ 11.313 22.625
300 $3.2410\times 10^{-7}$ 18.366 36.732
Table 2. Error in the estimate of the first non-zero eigenvalue for the sphere (randomly generated points on the sphere, $25^{3}$ Fourier basis functions on $\Omega=\left(-2,2\right)^{3}$, $d_{n}=\left(\exp\left(4\sqrt{2\pi/T}\right)+\exp\left(4\sqrt{\left\|\boldsymbol{\omega}_{n}\right\|_{2}}\right)\right)^{2}$, $T=4$).
Table 2 again shows high-order convergence. Overall, this method of finding eigenvalues seems to be more reliable and theoretically justified than attempting to produce a differentiation matrix from interpolation followed by differentiation of the basis functions, which can produce incorrect results, even for small $\lambda$ for a variety of collocation methods (see the discussion at the end of Subsection 2.2 in [yan23]). Its use is also not limited to eigenvalues; any solvability problem relying on parameters can be investigated similarly.
5. Conclusions
We successfully developed, analyzed, and tested methods for Hermite-Birkhoff interpolation and PDEs on surfaces. The methods rely on underdetermined Fourier extensions, and work on unstructured point clouds with very few conditions for convergence. We have shown that our method produces solutions that approximately solve PDEs of interest in an extremely general setup (Propositions 9 and 10); convergence rates could then be estimated for a wide range of PDEs. Convergence is also shown in a general setting without rates by Proposition 12.
Of particular interest, our method works for surface PDEs on unstructured point clouds and is able to achieve arbitrarily high rates of convergence. For bulk problems, our method could be seen as an alteration of earlier, symmetric Hermite RBF methods (see, for example, [sun94, frank98, liu23]). Hermite RBFs do not seem to be commonly applied to surface PDEs outside of one approach for RBF finite differences [shaw19]. The setup of traditional Hermite RBF approaches, however, can be quite difficult, and may help explain why such methods are not widespread compared to non-symmetric RBF methods. We also find that it is typically easier to work with Fourier series directly to set up the correct linear system rather than with the often unwieldy functions that arise with Hermite RBFs.
Furthermore, our method allows for more stable methods of solution; solving the optimization problem by complete orthogonal decomposition or singular value decomposition tends to produce superior results compared to actually forming $\boldsymbol{\Phi}_{N_{b}}$ when the system is ill-conditioned. Again, this is related to the fact that the condition number of $\boldsymbol{\Phi}_{N_{b}}=\boldsymbol{V}_{N_{b}}\boldsymbol{V}_{N_{b}}^{*}$ is the square of the condition number of $\boldsymbol{V}_{N_{b}}$; forming $\boldsymbol{\Phi}_{N_{b}}$ is quite similar to using Hermite RBFs, but with positive definite functions on a box rather than all of $\mathbb{R}^{m}$.
Our future work seeks to take advantage of the flexibility of the method for problems that would be intractable using more traditional approaches. Spacetime methods have shown promise in numerical testing, as have methods for producing conformal parametrizations of surfaces defined solely by point clouds. Particularly, such as in the case of conformal mapping, the method allows for an accurate solution to be found to problems with many possible solutions. A similar approach to Subsection 4.2 may also allow a range of solvability problems to be investigated numerically in a rigorously justified manner. Overall, we expect that our method could be useful for a large number of problems since its setup is nearly universal for linear PDEs (on surfaces), and since it is meshfree and high-order.
Acknowledgments
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), [funding reference number RGPIN 2022-03302].
[…]